a patrol distributor records shows that sales average 3500 its per day with a stadard deviation 250 .. assume that the sales are normally distributed, find the probability of daily sales between 3300 litre 3800 litre?
1
Expert's answer
2020-09-09T18:50:45-0400

Distribution function is "P(X\le x)=F(x)=\frac{1}{250\sqrt{2\pi}}\int\limits_{-\infty}^x e^{-\frac{(t-3500)^2}{2\cdot 250^2}}dt"

If we take "u=\frac{t-3500}{250}", then "\frac{(t-3500)^2}{2\cdot 250^2}=\frac{u^2}{2}" and "dt=250du", so "F(x)=\frac{1}{250\sqrt{2\pi}}\int\limits_{-\infty}^\frac{x-3500}{250} e^{-\frac{u^2}{2}}\cdot 250du=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^\frac{x-3500}{250} e^{-\frac{u^2}{2}}du="

"=\Phi(\frac{x-3500}{250})", where "\Phi(y)=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^y e^{-\frac{u^2}{2}}du"

So the required probability is "P(3300\le X\le 3800)=F(3800)-F(3300)="

"=\Phi(1.2)-\Phi(-0.8)"

From the table of standard normal distribution https://www.math.arizona.edu/~rsims/ma464/standardnormaltable.pdf we can see that "\Phi(1.2)\approx 0.88493" and "\Phi(-0.8)\approx 0.21186", so "P(3300\le X\le 3800)\approx0.88493-0.21186=0.67307"

Answer: "0.67307"